# Determinant Estimation under Memory Constraints and Neural Scaling Laws

### v1

@@ Abstract
Calculating or accurately estimating log-determinants of large positive semi-definite matrices is of fundamental importance in many machine learning tasks. While its cubic computational complexity can already be prohibitive, in modern applications, even storing the matrices themselves can pose a memory bottleneck. To address this, we derive a novel hierarchical algorithm based on block-wise computation of the LDL decomposition for large-scale log-determinant calculation in memory-constrained settings. In extreme cases where matrices are highly ill-conditioned, accurately computing the full matrix itself may be infeasible. This is particularly relevant when considering kernel matrices at scale, including the empirical Neural Tangent Kernel (NTK) of neural networks trained on large datasets. Under the assumption of neural scaling laws in the test error, we show that the ratio of pseudo-determinants satisfies a power-law relationship, allowing us to derive corresponding scaling laws. This enables accurate estimation of NTK log-determinants from a tiny fraction of the full dataset; in our experiments, this results in a $\sim$100,000$\times$ speedup with improved accuracy over competing approximations. Using these techniques, we successfully estimate log-determinants for dense matrices of extreme sizes, which were previously deemed intractable and inaccessible due to their enormous scale and computational demands.
@@ 1 Introduction
Recently, the empirical Neural Tangent Kernel (NTK) has become a prominent theoretical and practical tool for studying the behavior of neural networks both during training and at inference (Novak et al., , 2022). The Gram matrix associated with the NTK has been used in lazy training (Chizat et al., , 2019) and shows promise as a tool to obtain uncertainty quantification estimates (Immer et al., , 2021). Its log-determinant—highly sensitive to small eigenvalues—has also received recent attention, both as a quantity of interest in model selection techniques that rely upon marginal likelihood approximation (Immer et al., , 2023; Hodgkinson et al., 2023b, ) and as a way to quantify the complexity of a learning problem (Vakili et al., , 2022). Similarly, both the NTK log-determinant and closely related quantities have recently appeared in quantification of generalization error via PAC-Bayes bounds (Hodgkinson et al., 2023c, ; Kim et al., , 2023). While several approximations have been proposed (Mohamadi et al., , 2023), convergence has only been shown in spectral norm, so they can only be expected to capture the overall behavior of the full NTK.
In practice, computing the log-determinant of the empirical NTK’s Gram matrix is a formidable task: besides suffering from the pathologies that ill-conditioned matrices are subject to at scale, the NTK appears to have its own peculiarities that make the problem particularly challenging. Indeed, the task is considered so impenetrable as to be universally avoided, since the NTK corresponding to most relevant datasets cannot be stored in memory. For example, storing the NTK corresponding to the relatively small MNIST dataset requires 2.9 terabytes of memory. The equivalent object for ImageNet-1k requires 13.1 exabytes, an order of magnitude larger than CERN’s current data storage capacity (Smith, , 2023). While it is appealing to consider only a small subset of the training dataset when considering such approximations, these estimates for the log-determinant can incur significant bias. Furthermore, we will find that leading approaches for dealing with log-determinants of empirical NTK Gram matrices using sketching and other Monte Carlo approximations tend to be highly inaccurate.
@@
The central contributions of this work address the issues faced when computing log-determinants of large matrices. We are primarily interested in matrices that cannot fit into memory. To compute an accurate baseline, we derive MEMDET, a memory-constrained algorithm for determinant computation. This facilitates exact calculation of log-determinants of NTK Gram matrices corresponding to neural networks with several million parameters, for the first time on datasets this large. We then provide a novel approximation technique based on the scaling behavior of a wide class of kernel matrices, by appealing to neural scaling laws. Scaling laws have played a prominent role in machine learning theory and practice, providing insight into the asymptotic behavior of generalization error (Li et al., , 2023; Vakili et al., , 2022), as well as guidance for compute-optimal resource allocation when deploying deep learning models at scale (Kaplan et al., , 2020; Hoffmann et al., , 2022). Specifically,
@@
  • we provide a high-performance Python package detkit [^1] For details see Appendix I., which implements the presented algorithms and can be used to reproduce the results of this paper.
Crucially, we demonstrate that our approximation technique is able to obtain estimates with lower error than incurred by reducing the numerical precision used in explicit computation. Our approximation techniques render an impractical task virtually routine, as shown in Table 1. To the best of our knowledge, this is the first time that the full empirical NTK corresponding to a dataset of this scale has been computed.
Table 1: Comparison of stochastic Lanczos quadrature (SLQ, with degree $l$, $s$ Monte Carlo samples, and full re-orthogonalization), MEMDET (Algorithm 3), and FLODANCE (Algorithm 1) on a dense $500000$\times$500000$$ NTK matrix for a ResNet50 model trained on CIFAR-10 with $50000$ datapoints. MEMDET computes the exact log-determinant and serves as the benchmark, with relative errors of other methods measured against it. Costs and wall time are based on an NVIDIA H100 GPU (\$2/hour) and an 8-core 3.6GHz CPU (\$0.2/hour) using Amazon pricing.
@@
MEMDET LDL, $n_{b}=32$ 41,667 0% \$364 8.8 days
FLODANCE $n_{s}=500,q=0$ 0.04 4% \$0.04 1 min
@@ 2 Computing Determinants at Scale
The computation of determinants of large matrices, particularly those expected to be highly ill-conditioned, is widely considered to be a computationally “ugly” problem, which should be avoided, if possible (Axler, , 1995). However, this is not always an option, for example when the determinant represents a volume form, is required in a determinantal point process, or plays a role in training GPs via empirical Bayes. Although certain limiting behaviors under expectation are well characterized (Hodgkinson et al., 2023a, ), they provide only a coarse approximation in finite-sample settings. The quadratic memory cost and cubic computational complexity of naïve implementations make exact computation prohibitive, necessitating the use of lower-precision or randomized methods at scale. While approximate methods can be useful in their own right, exact computation remains essential, at the very least, to establish meaningful baselines. In our experiments, due to the pathological spectral behavior of the matrices we consider, we will see that our proposed exact method is comparable in speed to the state-of-the-art Monte Carlo approximation techniques.
@@ 2.1 Low-Precision Arithmetic
Among the most common techniques for circumventing memory and computational bottlenecks when dealing with with large-scale calculations in numerical linear algebra is to cast numerical values into a lower precision, usually 32-bit, 16-bit, or even 8-bit floating point values, instead of 64-bit (double precision) values that would otherwise be used.
@@ 2.2 MEMDET: A Memory-Constrained Algorithm for Log-Determinant Computation
Conventional methods for computing the determinant of matrices include LU decomposition (for generic matrices), LDL decomposition (for symmetric matrices), and Cholesky decomposition (for symmetric positive-definite matrices). These methods also simultaneously provide the determinants of all sub-matrices $\mathbf{M}_{[:k,:k]}$, consisting of the first $k$ rows and columns ($k=1,\dots,m$) of the $m\times m$ matrix $\mathbf{M}$. This will be beneficial in our applications.
The LU decomposition has a computational complexity of approximately $\frac{2}{3}m^{3}$, while both LDL and Cholesky decompositions have a complexity of approximately $\frac{1}{3}m^{3}$. While computational complexity is a primary concern for large-scale determinant computation, memory limitation poses an even greater challenge. These methods require substantial memory allocation, either the size of the array if written in-place, or twice the size if the input array is preserved.
To address this, we present MEMDET, a memory-constrained algorithm for large-scale determinant computation. Below is a sketch of the algorithm, with details of its implementation, memory and computational aspects provided in Appendix C, and a description of the software implementing this algorithm provided in Appendix I. We focus on the algorithm for generic matrices using LU decomposition, though the LDL and Cholesky decompositions follow similarly.
Consider the $2\times 2$ block LU decomposition of $\mathbf{M}$ (see for instance Dongarra et al., 1998b (, Chapter 5.4)) with the first block, $\mathbf{M}_{11}$, of size $b\times b$, $b<m$, as
@@
[…]
Figure 1: Schematic diagram of MEMDET illustrating the efficient block processing order for LU decomposition (Algorithm 2, left panel) and LDL/Cholesky decompositions (Algorithm 3, Algorithm 4, right panel). The detailed ordering strategy is described in Appendix D.
@@
The computational cost of this procedure remains the same, independent of the number of blocks, $n_{b}$ (see Section E.1). However, increasing the number of blocks reduces memory usage at the expense of higher data transfer between disk and memory. Efficient implementation minimizes this by processing blocks in an order that reduces the loading of $\mathbf{B}$ and $\mathbf{C}$. The left panel of Figure 1 shows one such order, illustrating the procedure at the $k$-th iteration. Processing the blocks $\mathbf{M}_{ij}$ starts from the last row of the matrix and moves upward. During the horizontal and vertical traverses on the path shown in the figure, the memory blocks $\mathbf{C}$ and $\mathbf{B}$ can remain in memory, avoiding unnecessary reloading. Once the procedure reaches the block $(k+1,k+1)$, the memory block $\mathbf{S}$ can be directly read into $\mathbf{A}$ instead of being stored on disk as $\mathbf{M}_{k+1,k+1}$, initiating the block $\mathbf{A}$ for the next iteration. Consequently, only the blocks shown in blue need to be stored in the scratchpad space, while those in dark blue are already stored in the current state of the algorithm. This algorithm can be further modified to eliminate the need for $\mathbf{A}$ and use the memory space of $\mathbf{S}$ instead, but this would require storing the additional gray blocks on the scratchpad space in addition to the blue blocks. A pseudo code and further efficient implementation details of the presented method can be found as Algorithm 2 in Appendix D.
@@
The algorithms for LDL and Cholesky decompositions follow a similar procedure with necessary adjustments for symmetric and symmetric positive-definite matrices, as detailed in Algorithm 3 and Algorithm 4. These modifications include processing only the lower (or upper) triangular part of the matrix and handling permutations and diagonal scaling in the LDL decomposition. The right panel of Figure 1 shows one possible block ordering for the algorithm for LDL and Cholesky decomposition: the ordering is optimal in minimizing the number of reads and writes, but is not unique. Other block processing orderings with the same amount of data access also exist, and these will be further discussed in Appendix D. A detailed analysis of the computational complexity and memory/disk data transfers of the algorithm is provided in Appendix E. Implementation of MEMDET algorithm can be found in LABEL:list:detkit1.
@@
Although the Neural Tangent Kernel (NTK) matrices we work with are symmetric and positive definite (SPD), we do not solely rely on Cholesky decomposition, despite it being the most suitable method for SPD matrices. This is because NTK matrices can lose their positive-definiteness with even the smallest perturbations causing small eigenvalues to become negative, such as when converting from 64-bit to 32-bit precision for efficient computation. As a result, the Cholesky decomposition becomes unstable and fails, necessitating the use of LU and LDL decompositions, suitable for more general matrices. We note that the block computations of LU decomposition can become unstable as the matrices deviate from symmetric and positive-definiteness (Demmel et al., , 1995), requiring pivoting of the blocks. However, in our empirical study of NTK matrices, we found that LU decomposition works well without block pivoting, though we do consider pivoting within each block.
@@ 3 Scaling Law for the Determinant
The NTK was first derived in the context of neural networks. However, the quantity is well defined for more a general class of functions. For a continuously differentiable function $f_{\theta}:\mathcal{X}\to\mathbb{R}^{d}$, we define its (empirical) NTK to be
$\kappa_{\theta}(x,x^{\prime})\coloneqq\mathbf{J}_{\theta}\big{(}f_{\theta}(x)\big{)}\mathbf{J}_{\theta}\big{(}f_{\theta}(x^{\prime})\big{)}^{\intercal},$ (2)
where $\mathbf{J}_{\theta}(f_{\theta}(x))\in\mathbb{R}^{d\times p}$ is the Jacobian of the function $f_{\theta}$ with respect to the flattened vector of its parameters $\theta\in\mathbb{R}^{p}$, evaluated at the point $x$. Note that the assumption of continuous differentiability is often relaxed in practice, and we have $\kappa_{\theta}(x,x^{\prime})\in\mathbb{R}^{d\times d}$. When computing the NTK across $n_{1}$ and $n_{2}$ datapoints, it is typical to arrange the $n_{1}n_{2}$ matrices as an $n_{1}d\times n_{2}d$ matrix. The scaling laws we derive in the sequel hold for a class of kernel functions satisfying certain decay properties on the eigenvalues of an associated integral operator (see Appendix F for a precise statement, taken from Li et al., (2023)). We remark that this class of kernels includes NTKs, as demonstrated in Bietti & Mairal, (2019); Bietti & Bach, (2021); Lai et al., (2023).

### v2

@@ Abstract
Calculating or accurately estimating log-determinants of large positive definite matrices is of fundamental importance in many machine learning tasks. While its cubic computational complexity can already be prohibitive, in modern applications, even storing the matrices themselves can pose a memory bottleneck. To address this, we derive a novel hierarchical algorithm based on block-wise computation of the LDL decomposition for large-scale log-determinant calculation in memory-constrained settings. In extreme cases where matrices are highly ill-conditioned, accurately computing the full matrix itself may be infeasible. This is particularly relevant when considering kernel matrices at scale, including the empirical Neural Tangent Kernel (NTK) of neural networks trained on large datasets. Under the assumption of neural scaling laws in the test error, we show that the ratio of pseudo-determinants satisfies a power-law relationship, allowing us to derive corresponding scaling laws. This enables accurate estimation of NTK log-determinants from a tiny fraction of the full dataset; in our experiments, this results in a $\sim$100,000$\times$ speedup with improved accuracy over competing approximations. Using these techniques, we successfully estimate log-determinants for dense matrices of extreme sizes, which were previously deemed intractable and inaccessible due to their enormous scale and computational demands.
@@ 1 Introduction
Recently, the empirical Neural Tangent Kernel (NTK) has become a prominent theoretical and practical tool for studying the behavior of neural networks both during training and at inference (Novak et al., , 2022; Hodgkinson et al., , 2025). The Gram matrix associated with the NTK has been used in lazy training (Chizat et al., , 2019) and shows promise as a tool to obtain uncertainty quantification estimates (Immer et al., , 2021; Wilson et al., , 2025). Its log-determinant—highly sensitive to small eigenvalues—has also received recent attention, both as a quantity of interest in model selection techniques that rely upon marginal likelihood approximation (Immer et al., , 2023; Hodgkinson et al., 2023b, ) and as a way to quantify the complexity of a learning problem (Vakili et al., , 2021). Similarly, both the NTK log-determinant and closely related quantities have recently appeared in quantification of generalization error via PAC-Bayes bounds (Hodgkinson et al., 2023c, ; Kim et al., , 2023). While several approximations have been proposed (Mohamadi et al., , 2023), convergence has only been shown in spectral norm, so they cannot be expected to capture the overall behavior of the full NTK.
In practice, computing the log-determinant of the empirical NTK’s Gram matrix is a formidable task: besides suffering from the pathologies that ill-conditioned matrices are subject to at scale, the NTK appears to have its own peculiarities that make the problem particularly challenging. Indeed, the task is considered so impenetrable as to be universally avoided, since the NTK corresponding to most relevant datasets cannot be stored in memory. For example, storing the NTK corresponding to the relatively small MNIST dataset requires 2.9 terabytes of memory. The equivalent object for ImageNet-1k requires 13.1 exabytes, an order of magnitude larger than CERN’s current data storage capacity (Smith, , 2023). While it is appealing to consider only a small subset of the training dataset when making necessary approximations, naïve estimates for the log-determinant can incur significant bias. Furthermore, we will find that leading approaches for dealing with log-determinants of empirical NTK Gram matrices using sketching and other Monte Carlo approximations tend to be highly inaccurate.
Scaling laws have played a prominent role in machine learning theory and practice, providing insight into the asymptotic behavior of generalization error (Li et al., , 2023; Vakili et al., , 2021), as well as guidance for compute-optimal resource allocation when deploying deep learning models at scale (Kaplan et al., , 2020; Hoffmann et al., , 2022). It turns out that for large kernel matrices, the ratio of successive determinants can be cast in terms of the error of an associated Gaussian process. This enables known scaling laws to be deployed in estimating the log-determinant, with surprising accuracy. However in order to measure this accuracy against a reliable baseline, a memory-constrained method for exact out-of-core log-determinant computation is also required, and is of independent interest.
@@
The central contributions of this work address the issues faced when computing log-determinants of large matrices. We are primarily interested in matrices that cannot fit into memory. To compute an accurate baseline, we derive MEMDET, a memory-constrained algorithm for determinant computation. This facilitates exact calculation of log-determinants of NTK Gram matrices corresponding to neural networks with several million parameters. We then provide a novel approximation technique, FLODANCE, based on the scaling behavior of a wide class of kernel matrices, by appealing to neural scaling laws. In detail, our main contributions are:
@@
  • we provide a high-performance Python package detkit, which implements the presented algorithms and can be used to reproduce the results of this paper.
Crucially, we demonstrate that our approximation technique is able to obtain estimates with lower error than incurred by reducing the numerical precision in explicit computation. Our approximation techniques render an impractical task virtually routine, as shown in Table 1. To the best of our knowledge, this is the first time that the full empirical NTK corresponding to a dataset of this scale has been computed.
Table 1: Comparison of stochastic Lanczos quadrature (SLQ, with degree $l$, $s$ Monte Carlo samples, and full re-orthogonalization), MEMDET (Algorithm 3), and FLODANCE (Algorithm 1) on a dense $500000$\times$500000$$ NTK matrix for a ResNet50 model trained on CIFAR-10 with $50000$ datapoints. MEMDET computes the exact log-determinant and serves as the benchmark, with relative errors of other methods measured against it. Costs and wall time are based on an NVIDIA H100 GPU (\$2/hour) and an 8-core 3.6GHz CPU (\$0.2/hour) using Amazon pricing and include NTK formation from a pre-trained network.
@@
MEMDET LDL, $n_{b}=32$ $41667$ 0% \$601 13.8 days
FLODANCE $n_{s}=500$,0 $q=0$ 0.04 4% \$0.04 1 min
@@
A summary of the notation used throughout the paper is provided in Appendix A. An overview of related work in linear algebra is given in Appendix B. Computational challenges when dealing with NTK matrices are then discussed in Appendix C. Implementation and performance analysis of MEMDET are provided in Appendices D and E. Required background for neural scaling laws, proofs of our theoretical results, and further analysis of FLODANCE appear in Appendices F and G. Comparison of various log-determinant methods is given in Appendix H. Finally, an implementation guide for detkit appears as Appendix I.
@@ 2 Computing Determinants at Scale
The computation of determinants of large matrices, particularly those expected to be highly ill-conditioned, is widely considered to be a computationally “ugly” problem, which should be avoided wherever possible (Axler, , 1995). However, this is not always an option, e.g., when the determinant represents a volume form, is required in a determinantal point process, or plays a role in training GPs via empirical Bayes. Although certain limiting behaviors under expectation are well characterized (Hodgkinson et al., 2023a, ), they provide only a coarse approximation in finite-sample settings. The quadratic memory cost and cubic computational complexity of naïve implementations make exact computation prohibitive, necessitating the use of lower-precision or randomized methods at scale. While approximate methods can be useful in their own right, exact computation remains essential, at the very least to establish meaningful baselines. In our experiments, due to the pathological spectral behavior of the matrices we consider, we will see that our proposed exact method is comparable in speed to the state-of-the-art Monte Carlo approximation techniques.
@@ 2.1 Low-Precision Arithmetic
Among the most common techniques for circumventing memory and computational bottlenecks when dealing with large-scale calculations in numerical linear algebra is to cast numerical values into a lower precision, usually 32-bit, 16-bit, or even 8-bit floating point values, instead of 64-bit (double precision) values that would otherwise be used.
@@ 2.2 MEMDET: A Memory-Constrained Algorithm for Log-Determinant Computation
Conventional methods for computing the determinant of matrices include LU decomposition (for generic matrices), LDL decomposition (for symmetric matrices), and Cholesky decomposition (for symmetric positive-definite matrices). These methods also simultaneously provide the determinants of all leading principal submatrices $\mathbf{M}_{[:k,:k]}$ (for $k=1,\dots,m$), possibly after permuting $\mathbf{M}$ depending on the decomposition. This is beneficial in our applications.
The LU decomposition has a computational complexity of approximately $\frac{2}{3}m^{3}$, while both LDL and Cholesky decompositions have a complexity of approximately $\frac{1}{3}m^{3}$. While computational complexity is a concern for large-scale determinant computation, memory limitation poses an even greater challenge. These methods require substantial memory allocation, either the size of the array if written in-place, or twice the size if the input array is preserved.
To address this, we present MEMDET, a memory-constrained algorithm for large-scale determinant computation. Below is a sketch of the algorithm, with details of its implementation, memory and computational aspects provided in Appendix D, and a description of the software implementing this algorithm provided in Appendix I. We focus on the algorithm for generic matrices using LU decomposition, though the LDL and Cholesky decompositions follow similarly.
[…]
Figure 1: Schematic diagram of MEMDET illustrating the efficient block processing order for LU decomposition (Algorithm 2, left panel) and LDL/Cholesky decompositions (Algorithm 3, Algorithm 4, right panel). The detailed ordering strategy is described in Section D.1.
Consider the $2\times 2$ block LU decomposition of $\mathbf{M}$ (see for instance, Dongarra et al., 1998b (, Chapter 5.4)) with the first block, $\mathbf{M}_{11}$, of size $b\times b$, $b<m$, as
@@
The computational cost of this procedure remains the same, independent of the number of blocks, $n_{b}$ (see Section E.1). However, increasing the number of blocks reduces memory usage at the expense of higher data transfer between disk and memory. Efficient implementation minimizes this by processing blocks in an order that reduces the loading of $\mathbf{B}$ and $\mathbf{C}$. The left panel of Figure 1 shows one such order, illustrating the procedure at the $k$-th iteration. Processing the blocks $\mathbf{M}_{ij}$ starts from the last row of the matrix and moves upward. During the horizontal and vertical traverses on the path shown in the figure, the memory blocks $\mathbf{C}$ and $\mathbf{B}$ can remain in memory, avoiding unnecessary reloading. Once the procedure reaches the block $(k+1,k+1)$, the memory block $\mathbf{S}$ can be directly read into $\mathbf{A}$ instead of being stored on disk as $\mathbf{M}_{k+1,k+1}$, initiating the block $\mathbf{A}$ for the next iteration. Consequently, only the blocks shown in blue need to be stored in the scratchpad space, while those in dark blue are already stored in the current state of the algorithm. This algorithm can be further modified to eliminate the need for $\mathbf{A}$ and use the memory space of $\mathbf{S}$ instead, but this would require storing the additional gray blocks on the scratchpad space in addition to the blue blocks. A pseudo code and further efficient implementation details of the presented method can be found as Algorithm 2.
@@
The algorithms for LDL and Cholesky decompositions follow a similar procedure with necessary adjustments for symmetric and symmetric positive-definite matrices, as detailed in Algorithm 3 and Algorithm 4. These modifications include processing only the lower (or upper) triangular part of the matrix and handling permutations and diagonal scaling in the LDL decomposition. The right panel of Figure 1 shows one possible block ordering for the algorithm for LDL and Cholesky decomposition: the ordering is optimal in minimizing the number of reads and writes, but it is not unique. Other block processing orderings with the same amount of data access also exist, and these will be further discussed in Section D.1. A detailed analysis of the computational complexity and memory/disk data transfers of the algorithm is provided in Appendix E. Implementation of MEMDET algorithm can be found in LABEL:list:detkit1.
@@
Although the NTK matrices we work with are symmetric and positive definite (SPD), we do not solely rely on Cholesky decomposition, despite it being the most suitable method for SPD matrices. This is because NTK matrices can lose their positive-definiteness with even the smallest perturbations causing small eigenvalues to become negative, such as when converting from 64-bit to 32-bit precision for efficient computation. As a result, the Cholesky decomposition becomes unstable and fails, necessitating the use of LU and LDL decompositions, suitable for more general matrices. We note that the block computations of LU decomposition can become unstable as the matrices deviate from symmetry and positive-definiteness (Demmel et al., , 1995), requiring pivoting of the blocks. However, in our empirical study of NTK matrices, we found that LU decomposition works well without block pivoting, though we do consider pivoting within each block.
@@ 3 Scaling Law for the Determinant
The NTK was first derived in the context of neural networks. However, the quantity is well defined for a more general class of functions. For a continuously differentiable function $f_{\boldsymbol{\theta}}:\mathcal{X}\to\mathbb{R}^{d}$, where $d$ is the dimensionality of the model’s output (e.g., the number of labels in a classification task), we define its (empirical) NTK as
$\kappa_{\boldsymbol{\theta}}(x,x^{\prime})\coloneqq\mathbf{J}_{\boldsymbol{\theta}}\big{(}f_{\boldsymbol{\theta}}(x)\big{)}\mathbf{J}_{\boldsymbol{\theta}}\big{(}f_{\boldsymbol{\theta}}(x^{\prime})\big{)}^{\intercal},$ (2)
where $\mathbf{J}_{\boldsymbol{\theta}}(f_{\boldsymbol{\theta}}(x))\in\mathbb{R}^{d\times p}$ is the Jacobian of the function $f_{\boldsymbol{\theta}}$ with respect to the flattened vector of its parameters $\boldsymbol{\theta}\in\mathbb{R}^{p}$, evaluated at the point $x$. The assumption of continuous differentiability is often relaxed in practice.
Note that $\kappa_{\boldsymbol{\theta}}(x,x^{\prime})\in\mathbb{R}^{d\times d}$, so computing the NTK across $n$ datapoints yields a 4th-order tensor of shape $(n,n,d,d)$. For computational purposes, this is typically flattened into a two-dimensional block matrix of size $nd\times nd$, where each $(i,j)$-block corresponds to the $d\times d$ matrix $\kappa(x_{i},x_{j})$.
The scaling laws we derive apply to kernel families satisfying decay conditions on the eigenvalues of an associated integral operator (see Section F.1 for a precise statement drawn from Li et al., (2023)). This class of kernels includes NTKs, as shown by Bietti & Mairal, (2019); Bietti & Bach, (2021); Lai et al., (2023); see Figure 2.